OpenStudy (anonymous):

write a quadratic equation having the given numbers as solutions. -6 and -4

OpenStudy (perl):

you can work backwards The first solution gives us: x = -6 x+6 = 0 The second solution gives us: x = -4 x + 4= 0 now combine them (x+6)(x+4) =0 expand the left side

OpenStudy (perl):

(x+6)(x+4) =0 to expand the left side we can use the distributive rule a * (b + c) = a*b + a*c In our problem treat 'x + 6' as 'a' (x+6)*(x + 4) = (x+6)*x + (x+6)*4

OpenStudy (perl):

then use distribute law again

OpenStudy (perl):

(x+6)*(x + 4) = (x+6)*x + (x+6)*4 = x*x + 6*x + x*4 + 6*4